OpenStudy (anonymous):

Find the solution set for the open sentence with the given replacement set. 2x + 7 = 19, {4, 5, 6, 7}

OpenStudy (loinducoeur):

If you're trying to find x, then you can do these following steps to get x:\[2x+7=19\]\[2x=19-7\]\[2x=12\]\[x=12\div2\] And I'm sure you can do that last sum to get the final answer (x). :)